OpenStudy (anonymous):

How would I solve this. (The ----- means a fraction bar) (x+16)=3x ------ 3

OpenStudy (anonymous):

\[\frac{ x+16 }{ 3 }=3x\] ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

too many pppl here now. good luck

OpenStudy (anonymous):

Well.... you would multiply by 3 on both sides. subtract the x and there you go?

OpenStudy (anonymous):

x+16 = 9x 16=8x

OpenStudy (anonymous):

That should be simplified enough